# RomanSetu: Efficiently unlocking multilingual capabilities of Large Language Models via Romanization

### v2

@@ 4.5 Evaluation Metrics
For other NLG tasks, we report Rouge-L Lin (2004) scores. We primarily report evaluations on the native script outputs. In the case of romanized outputs, we transliterate them back to native script using IndicXlit combined with a unigram word-level language model Madhani et al. (2023). Combining word-level results with a unigram language model is efficient and preserves most gains of incorporating a complete n-gram language model Kirov et al. (2024). In this native script evaluation, the roman script models outputs can also have transliteration errors, which put them at a disadvantage compared to native script model. To purely compare task performance, we also report results on romanized outputs. For this, we romanize the native script outputs using IndicXlit. Since the final output is an artifact of the same romanization model, this evaluation compares only the task performance.
4.6 Models Compared
@@ 4.7 Prompting and Decoding
For the base and CPT models, we do 1-shot and 3-shot prompting. For IFT models, we experiment with 0-shot prompting. The prompt templates for various tasks are shown in Appendix B. For generation tasks, we use greedy decoding. For classification tasks, we completed the prompts with various possible options as per class labels and compute the output probabilities. We choose the class that maximizes the probability (Liu et al., 2021).
@@ 5 Results
In this section, we present results of various experiments comparing romanized and native representations and discuss the results.
We aim to address the following research questions:
@@
Table 1: Fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2022) for native (N), romanized with IndicXLIT (R) and romanized with IndicNLP library (R-IndicNLP). Bold indicates lowest fertility.
Efficiency Gains with Romanization
To address this question, we compute fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2022). Fertility refers to the number of tokens per word generated by the Llama 2 tokenizer Touvron et al. (2023). We compare the native script against two forms of romanized text: transliteration using IndicXlit Madhani et al. (2023) and rule-based extended ITRANS transliteration scheme implemented in the IndicNLP library Kunchukuttan (2020). Table 1 presents the fertility scores on the Llama 2 model for the languages we experimented with. We observed that the fertility (and hence sequence length) in the native script is at least twice that of the IndicXlit romanized text. In languages like Gujarati, the native script feritility is very large (greater than 4 times the romanized fertility). Hence, we can expect significant efficiency gains when processing romanized text over native text in terms of reduced memory consumption, reduced generation time and increased maximum sequence length limit.
Furthermore, romanization using IndicNLP Kunchukuttan (2020) has slightly higher fertility compared to IndicXlit. While we have used IndicXlit-based transliterations in this work for reasons explained earlier, it would be interesting to explore deterministic and reversible romanizations like IndicNLP at the cost of slightly increased inefficiency. The advantage of reversible mapping would be avoiding transliteration errors when mapping the outputs back to native script.
@@
Cross-lingual Representation Similarity
@@
Machine Translation
We discuss results of machine translation, a canonical multilingual language generation task. Table 3 (Rows XX-En translation and En-XX translation) shows results on machine translation both in and out of English.
For translation into English, the base model translation quality is roughly equivalent for native and Roman scripts. We see that continual pre-training helps improve translation quality significantly for both representations. Romanized scripts show significantly more improvement on the CPT models compared to native scripts. We hypothesize that romanization is better able to align cross-lingual representations between English and other languages to improve translation performance. Romanized IFT models also perform better than their native counterparts, albeit the gap is now narrower compared to CPT models.
@@
Table 4 shows the effect of the the choice of script when evaluating translation out of English to each language. We also report evaluation on Roman script outputs for both models, and see that the romanized IFT models report even larger gains in this evaluation setting. This suggests that the true gains from romanization might be larger and it is not reflected in native script evaluation on account of errors in Roman to native transliteration that is performed for the evaluation.
@@
Table 3: Results on NLG Benchmarks - averaged across languages (RougeL score for XLSum and IndicHeadline; chrF for MT. En-$X$ refers to out-of-English and $X$-En refers to into-English directions. N and R indicate whether the model was trained and decoded on native or romanized script. Evaluation is always in native script.
@@
Language Generation Tasks
@@
[…]
@@
[…]
Language Understanding Tasks
@@
Knowledge and Reasoning tasks
Table 6 shows the results on translated versions of various knowledge and reasoning benchmarks such as MMLU (Hendrycks et al., 2021b), BoolQ (Clark et al., 2019b), ARC (Clark et al., 2018b) and CommonsenseQA (Talmor et al., 2019). For most of the tasks, the results of romanized script are competitive with native script. At the same time, they are efficient to process, as discussed earlier. For IndicQA without context (open domain question answering task), we observe that the romanized model shows ~10 F1 point gain over the native script model. This demonstrates the overall effectiveness of utilizing romanization for non-Latin script languages.
Discussion
From the analysis and results presented earlier, we can summarize the following. Romanized representations are significantly more efficient compared to native script representations. They are better aligned to English representations compared to native script representations, hence they are better for cross-lingual transfer. Across a range of tasks and languages, romanized representations provide competitive or better performance over their native script counterparts (detailed results per language and task are presented in Appendix D. Specifically, for a generation task like machine translation we see significant gains while translating out of English with romanized representations. In Appendix C, we show some examples of romanized and native script translations, illustrating the gains from romanization. These results indicate that romanization is a promising alternative to extending the capabilities of English LLMs to other languages using non-Roman scripts.
6 Conclusion and Future Work
@@ 7 Limitations
In this work, we chose to use natural romanizations since they tended to have lower fertility and might have been seen in the base LM pre-training. Transliterating the outputs back to native script can be lossy. We use a high-quality transliteration model to minimize the transliteration errors. However, such high quality transliterations might not be available for all languages. Using a deterministic and reversible transliteration scheme can alleviate this problem.
The experiments in this work have been conducted on Indian languages. Indian languages are representative of the scenarios we study and we cover multiple language families and script with varying resource status. Previous studies have also used Indic languages as a case study for multilinguality and multi-script scenarios Moosa et al. (2022); Khemchandani et al. (2021). Hence, we believe the findings should generalize to other language families - but further experimentation can confirm this hypothesis.
We have done only a limited amount of pretraining data in our experiments due to compute constraints. Resource constraints led us to limit our experiments to a 7B model, but experiments with larget models can yield more insights. Larger models and pre-training on larger datasets might benefit the multilingual models Kaplan et al. (2020); Fernandes et al. (2023). We expect the broad conclusions to be the same - that romanized models are more efficient and better/comparable in task performance to native-script models.
Some of the evaluations have been done on translated benchmarks. While this is not ideal, most low-resource languages testsets across diverse tasks. Hence, multiple works have relied on translated benchmarks to measure multilingual performance - these can give indicative trends on performance across languages. More efforts are needed to created multilingual benchmarks for diverse tasks.
@@ 8 Ethics Statement
We use romanization as a way to align English with other languages using non-Latin scripts. Given the current state of LLMs, this seems like a practical direction to extend capabilities of the best LLMs to other languages. The intention is not to supplant the use of a native script (which is widely adopted and has a rich literary tradition) with romanized script, but to use romanization as a way to efficiently bring the benefits of LLM technology to low-resource languages written in non-Latin scripts. Further advancements are needed to extend this line of research to improve native script performance efficiently.
@@ Appendix B Prompt Templates
Appendix C Translation Examples
This section shows some examples of translations using the romanized and native-script IFT models (See Figures 13, 14, 15 and 16). We can see in these selected examples that romanized representaton performs better.

### v3

@@ 4.5 Evaluation Metrics
For other NLG tasks, we report Rouge-L Lin (2004) scores. We primarily report evaluations on the native script outputs. In the case of romanized outputs, we transliterate them back to native script using IndicXlit combined with a unigram word-level language model Madhani et al. (2023). Combining word-level results with a unigram language model is efficient and preserves most gains of incorporating a complete n-gram language model Kirov et al. (2024). In this native script evaluation, the roman script model outputs can also have transliteration errors, which put them at a disadvantage compared to the native script model. To purely compare task performance, we also report results on romanized outputs. For this, we romanize the native script outputs using IndicXlit. Since the final output is an artifact of the same romanization model, this evaluation compares only the task performance.
4.6 Models
@@ 4.7 Prompting and Decoding
For the base and CPT models, we do 1-shot and 3-shot prompting. For IFT models, we experiment with 0-shot prompting. The prompt templates for various tasks are shown in Appendix B. For generation tasks, we use greedy decoding. For classification tasks, we completed the prompts with various possible options as per class labels and computed the output probabilities. We choose the class that maximizes the probability (Liu et al., 2021).
@@ 5 Results
In this section, we present the results of various experiments comparing romanized and native representations and discuss the results. We aim to address the following research questions:
@@
Table 1: Fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2024) for native (N), romanized with IndicXLIT (R) and romanized with IndicNLP library (R-IndicNLP). Bold indicates lowest fertility.
Efficiency Gains with Romanization:
To address this question, we compute fertility scores on the FLORES-200 dev set (Goyal et al., 2022a; Costa-jussà et al., 2024). Fertility refers to the number of tokens per word generated by the Llama 2 tokenizer Touvron et al. (2023). We compare the native script against two forms of romanized text: transliteration using IndicXlit Madhani et al. (2023) and rule-based extended ITRANS transliteration scheme implemented in the IndicNLP library Kunchukuttan (2020). Table 1 presents the fertility scores on the Llama 2 model for the languages we experimented with. We observed that the fertility (and hence sequence length) in the native script is at least twice that of the IndicXlit romanized text. In languages like Gujarati, the native script fertility is very large (greater than 4 times the romanized fertility). Hence, we can expect significant efficiency gains when processing romanized text over native text in terms of reduced memory consumption, reduced generation time and increased maximum sequence length limit.
Furthermore, romanization using IndicNLP Kunchukuttan (2020) has slightly higher fertility compared to IndicXlit. While we have used IndicXlit-based transliterations in this work for reasons explained earlier, it would be interesting to explore deterministic and reversible romanizations like IndicNLP at the cost of slightly increased inefficiency. The advantage of reversible mapping would be avoiding transliteration errors when mapping the outputs back to the native script.
@@
Cross-lingual Representation Similarity:
@@
Machine Translation:
We discuss the results of machine translation, a canonical multilingual language generation task. Table 3 (Rows XX-En translation and En-XX translation) shows results on machine translation both in and out of English.
For translation into English, the base model translation quality is roughly equivalent for both native and roman scripts. We see that continual pretraining helps improve translation quality significantly for both representations. Romanized scripts show significantly more improvement on the CPT models compared to native scripts. We hypothesize that romanization is better able to align cross-lingual representations between English and other languages to improve translation performance. Romanized IFT models also perform better than their native counterparts, albeit the gap is now narrower compared to CPT models.
@@
Table 4 shows the effect of the the choice of script when evaluating translation out of English to each language. We also report evaluation on Roman script outputs for both models and see that the romanized IFT models report even larger gains in this evaluation setting. This suggests that the true gains from romanization might be larger and it is not reflected in native script evaluation on account of errors in Roman to native transliteration that is performed for the evaluation.
@@
Table 3: Results on NLG Benchmarks - averaged across languages (RougeL score for XLSum and IndicHeadline; chrF for MT. En-$X$ refers to out-of-English and $X$-En refers to into-English directions. N and R indicate whether the model was trained and decoded on native or romanized script. Evaluation is always in the native script.
@@
Language Generation Tasks:
@@
IndicQA (with context) N 9.77 - 23.83 - 16.79
R 19.74 - 27.25 - 20.33
Table 5: Results on IndicXTREME NLU Testsets (average F1 score across languages). indicates 1-shot as scores for 0-shot were close to zero and non-interpretable so we made an exception for this task. N and R indicate whether the model was trained and decoded on native or romanized script. Evaluation is always in the native script.
@@
IndicQA (without context) N 1.48 - 5.56 - 1.13
R 12.96 - 17.62 - 16.87
Table 6: Results on Translated NLU Testsets (average Accuracy score across languages with the exception of IndicQA where use F1 score). indicates scores reported for 1-shot as scores for 0-shot were close to zero and non-interpretable so we made an exception for this task. N and R indicate whether the model was trained and decoded on native or romanized script. Evaluation is always in the native script.
Language Understanding Tasks:
@@
Knowledge and Reasoning tasks:
Table 6 shows the results on translated versions of various knowledge and reasoning benchmarks such as MMLU (Hendrycks et al., 2021b), BoolQ (Clark et al., 2019b), ARC (Clark et al., 2018b) and CommonsenseQA (Talmor et al., 2019). For most of the tasks, the results of the romanized script are competitive with the native script. At the same time, they are efficient to process, as discussed earlier. For IndicQA without context (open domain question answering task), we observe that the romanized model shows ~10 F1 point gain over the native script model. This demonstrates the overall effectiveness of utilizing romanization for non-Latin script languages.
Discussion:
From the analysis and results presented earlier, we can summarize the following. Romanized representations are significantly more efficient compared to native script representations. They are better aligned to English representations compared to native script representations, hence they are better for cross-lingual transfer. Across a range of tasks and languages, romanized representations provide competitive or better performance over their native script counterparts (detailed results per language and task are presented in Appendix D. Specifically, for a generation task like machine translation, we see significant gains while translating out of English with romanized representations. In Appendix C, we show some examples of romanized and native script translations, illustrating the gains from romanization. These results indicate that romanization is a promising alternative to extending the capabilities of English LLMs to other languages using non-Roman scripts.
We primarily base our study on English-only models due to their state-of-the-art performance across a diverse array of tasks and consistent improvements in capabilities. In contrast, multilingual models like BLOOM significantly lag behind English-only models, even on English tasks, and their progress has been slower. This disparity is due to the limited amount of English data and the lack of fine-tuning data for non-English languages used in training multilingual models. Therefore, improving English-only models is a practical way to enhance multilingual LLMs. We fine-tuned the BLOOM 7B model with the instruction dataset in both native and roman script. We observe that the roman script model outperforms the native script model, consistent with the findings on the Llama 2 base model. We provide the evaluation results for BLOOM 7B model across different tasks in Appendix E.
6 Conclusion
@@ 8 Limitations
In this work, we chose to use natural romanizations since they tended to have lower fertility and might have been seen in the base LM pretraining. Transliterating the outputs back to the native script can be lossy. We use a high-quality transliteration model to minimize the transliteration errors. However, such high-quality transliterations might not be available for all languages. Using a deterministic and reversible transliteration scheme can alleviate this problem.
The experiments in this work have been conducted on Indian languages. Indian languages are representative of the scenarios we study and we cover multiple language families and scripts with varying resource status. Previous studies have also used Indic languages as a case study for multilinguality and multi-script scenarios Moosa et al. (2023); Khemchandani et al. (2021). Hence, we believe the findings should generalize to other language families - but further experimentation can confirm this hypothesis.
We have done only a limited amount of pretraining data in our experiments due to compute constraints. Resource constraints led us to limit our experiments to a 7B model, but experiments with larger models can yield more insights. Larger models and pre-training on larger datasets might benefit the multilingual models Kaplan et al. (2020); Fernandes et al. (2023). We expect the broad conclusions to be the same - that romanized models are more efficient and better/comparable in task performance to native-script models.
Some of the evaluations have been done on translated benchmarks. While this is not ideal, most low-resource languages lack testsets across diverse tasks. Hence, multiple works have relied on translated benchmarks to measure multilingual performance - these can give indicative trends on performance across languages. More efforts are needed to create multilingual benchmarks for diverse tasks.
@@ 9 Ethics Statement
We use romanization as a way to align English with other languages using non-Latin scripts. Given the current state of LLMs, this seems like a practical direction to extend the capabilities of the best LLMs to other languages. The intention is not to supplant the use of a native script (which is widely adopted and has a rich literary tradition) with romanized script but to use romanization as a way to efficiently bring the benefits of LLM technology to low-resource languages written in non-Latin scripts. Further advancements are needed to extend this line of research to improve native script performance efficiently.
@@ Appendix B Prompt Templates
Appendix C Translation Examples
This section shows some examples of translations using the romanized and native-script IFT models (See Figures 13, 14, 15 and 16). We can see in these selected examples that romanized representation performs better.
Appendix D Language Specific Results
Appendix E Results on BLOOM 7B
We report evaluation results for 5 Indic languages for the tasks such as machine translation, headline generation and summarization on instruction fine-tuned BLOOM 7B in Tables 23, 24 and 25.
@@ Appendix E Results on BLOOM 7B
Model Shots Input text Hindi Gujarati Tamil Marathi Malayalam
IFT (BLOOM) - N 0 Native 51.03 48.39 52.1 45.63 49.94
IFT (BLOOM) - R 0 Romanized 53.74 50.95 50.09 48.06 48.93
Table 23: chrF scores for En-XX translation task on FLORES-200 devtest (Goyal et al., 2022a) across 5 Indic languages in both native and romanized script.
Model Shots Input text Hindi Gujarati Tamil Marathi Malayalam
IFT (BLOOM) - N 0-shot Native 7.44 4.36 7.66 0.8 3.7
IFT (BLOOM) - R 0-shot Romanized 20.96 15.87 25.4 13.46 19.42
Table 24: RougeL scores on IndicHeadline (Kumar et al., 2022) across 5 Indic languages in both native and romanized script.
Model Shots Input text Hindi Gujarati Tamil Marathi
IFT (BLOOM) - N 0-shot Native 1.73 2.84 2.56 2.73
IFT (BLOOM) - R 0-shot Romanized 17.16 10.74 8.63 9.36
Table 25: RougeL scores on XLSum (Hasan et al., 2021) across 5 Indic languages in both native and romanized script.